# The Belle II Collaboration

### v1

@@ V.1 Statistical model
The statistical model relies on the simultaneous measurement of event counts $n$ in two independent regions [^4] Referred to as channels in the pyhf documentation. – a signal region and a control region – using binned distributions of a given variable. These distributions are modelled as the sum of three different components (i.e. samples or templates, specified in Section V.3). Within the statistical model, a set of free and constrained parameters, $\bm{\eta}$ and $\bm{\chi}$ respectively, parametrize the variations of each template yield. The impact of each nuisance parameter $\chi$ is limited by a Gaussian constraint term $c_{\chi}$. The constraints are defined using a set of auxiliary measurements $\bm{a}$ (one measurement $a_{\chi}$ for each parameter), which correspond to component yield variations caused by various systematic effects. Thus, the statistical model $f$, in its most general form, can be written as
@@ V.2 Fitted regions and $B\rightarrow X_{c}\ell\nu$ shape correction
As detailed in Section IV.2, $B\rightarrow X_{c}\ell\nu$ decays appear to be poorly modelled. Normalization correction factors are calculated and applied before the fit whereas the correction of the $B\rightarrow X_{c}\ell\nu$ component shape is implemented directly into the fit. In order to constrain the shape of the $B\rightarrow X_{c}\ell\nu$ background in the signal region, distributions are fitted simultaneously in the signal region and CR0,low, the latter being orthogonal to the signal region and enriched in $B\rightarrow X_{c}\ell\nu$ events. This method not only corrects the $B\rightarrow X_{c}\ell\nu$ shape in the signal region, it also properly accounts for all correlations between correction factors and other fit parameters to determine a $B\rightarrow X_{u}\ell\nu$ branching fraction value that incorporates uncertainties related to the $B\rightarrow X_{c}\ell\nu$ background corrections. Using any control region to correct the shape of the $B\rightarrow X_{c}\ell\nu$ component relies on the assumption that the disagreement between data and simulation is similar in the signal region and the chosen control region. This assumption is supported by the fact that the ratio of $B\rightarrow X_{c}\ell\nu$ data over simulated data yields follows a similar trend in relevant kinematic variables in all three control regions as illustrated in Figure 6. To choose the control region used for correcting the signal region mismodeling, two considerations must be taken into account: the similarity in composition and shape of the $B\rightarrow X_{c}\ell\nu$ component in the control and signal region and the sample size relative to the signal region. Based on these two considerations, CR0,low is chosen for the fit.
@@ V.3 Fit templates
  • Signal ($B\rightarrow X_{u}\ell\nu$): this template includes all signal events whose generated kinematics fall in the chosen phase-space region (signal-in). Signal decays are modelled via the DFN framework [19]. The template is assigned an unconstrained normalization factor in the statistical model whose postfit value can be directly translated into a measurement of the branching fraction. This factor is fixed to its value from simulation in CR0,low where its contribution is negligible. An additional template (signal-out) is defined to cover the small fraction of signal events which pass the reconstruction-level selections given in Table 2 but whose generated kinematic parameters fall outside the target signal region (based on the generator-level selections). The resolution of the $E_{\ell}^{B}$ distribution is extremely good, allowing the $B\rightarrow X_{u}\ell\nu$-out yield to be neglected in fit 1. In fits 2 and 3, this component is kept as it represents 2% and 5% of all signal events in their respective signal regions. The template is assigned the same unconstrained normalization as the signal-in template but nuisance parameters can vary its normalization and shape independently.
  • $B\rightarrow X_{c}\ell\nu$ background: this template represents the largest source of background. It is composed of $B\rightarrow D\ell\nu$, $B\rightarrow D^{*}\ell\nu$, $B\rightarrow D^{**}\ell\nu$ and $B\rightarrow D^{**}_{\textrm{Gap}}\ell\nu$ decays. To capture the shape of this template in CR0,low, one free-floating normalization factor per bin, common to the signal and control-region distributions, is used. The normalization of this template is corrected prior to fitting, following the procedure described in Section IV.3. A constrained normalization parameter is assigned to the $B\rightarrow X_{c}\ell\nu$ template in each region to float the overall normalization within the uncertainties given in Table 3.
@@ V.5.1 $B\rightarrow X_{u}\ell\nu$ modeling
The resonant and non-resonant branching fractions (given in Table 1) are varied within their respective uncertainties therefore yielding one nuisance parameter per decay type. The modeling of $B\rightarrow X_{u}\ell\nu$ resonant modes relies on assumed form factor models which depend on a set of parameters extracted from lattice QCD and dedicated measurements. The form factor parameter variations are calculated using the eFFORT [46] package. The parameter values, uncertainties and correlation matrices are taken from the references given for each model in Section II. Two nuisance parameters are added to represent the uncertainties in $m_{b}^{\mathrm{KN}}=4.66\pm 0.04$ GeV and $a^{\mathrm{KN}}=1.3\pm 0.5$, including their covariance. In addition, to factor in differences between inclusive $B\rightarrow X_{u}\ell\nu$ models, a set of events is simulated via the Bosch, Lange, Neubert, Paz (BLNP) framework [10, 47]. The difference between the nominal fitted distribution and the BLNP distribution is added as a symmetric uncertainty. Moreover, applying a kaon veto to suppress charm background in the signal region also rejects $B\rightarrow X_{u}\ell\nu$ events where the $X_{u}$ system hadronizes as a pair of kaons. The branching fraction of $B\to K\bar{K}(X)\ell\nu$ decays (where $X$ is a charmless hadronic system) is not known and an uncertainty related to this contribution is therefore extracted by varying the production of $s$ quarks relative to $u$ or $d$ production, $\gamma_{s}$. The two values measured by the TASSO [48] ($\gamma_{s}=0.35\pm 0.05$) and JADE [49] ($\gamma_{s}=0.27\pm 0.06$) collaborations are averaged and the uncertainty is chosen to cover the full spread of $1\sigma$ variations from the two values, giving $\gamma_{s}=0.30\pm 0.09$ (which corresponds to a branching fraction of roughly 0.02%). Three sets of events with $\gamma_{s}$ equal to 0.21, 0.30 and 0.39 are produced. The sample simulated with $\gamma_{s}$ equal to its central value is used as nominal and the variations with respect to the other two distributions are added as systematic uncertainties. For each variation related to the $B\rightarrow X_{u}\ell\nu$ modeling, the hybrid weights are recalculated (see Equation 3).
@@ V.5.2 $B\rightarrow X_{c}\ell\nu$ modeling
All $B\rightarrow X_{c}\ell\nu$ and all $D$ meson decay branching fractions are varied within their respective uncertainties (see Table 1). The uncertainties of the unmeasured $B\to D^{(*)}\eta\ell\nu$ branching fractions are chosen to be equal to 100%. The branching fraction uncertainty of $B\to D_{0}^{*}\ell\nu$ decays is increased by about a factor of six to partially cover the large difference between the two measurements entering the HFLAV average calculation [50, 51]. The number of parameters related to the branching fractions of $D$ meson decays is truncated from 227 to one. For $B\rightarrow X_{c}\ell\nu$ decays, the form factor parameter variations are computed using the Hammer [52] package. The parameter values, uncertainties and correlation matrices are taken from the references given for each model in Section II.
@@ V.5.9 Continuum calibration
Simulated continuum events are corrected using the data-driven approach described in Section III.3. The full difference between the nominal distribution and the distribution corrected by the normalization factors is added as an uncertainty. Uncertainties on the continuum component shape also arise because of the limited size of the samples used to extract the correction factors. A total of 200 bootstrapped variations of the off-resonance experimental and simulated data samples are trained separately yielding as many sets of correction factors. The number of parameters implemented in the fit is truncated from 200 to five.
@@ V.5.11 Fit template normalizations
In fit 1, a normalization parameter is assigned to the other backgrounds template, which contributes to the total signal branching fraction uncertainty. The nuisance parameter is constrained by a Gaussian function with a width of 50% of the total component yields to vary the component normalization within a conservative range. In fits 2 and 3, the yields of non-semileptonic events being negligible, this parameter is not added but other nuisance parameters related to this template can vary both its normalization and shape. Furthermore, the control region and signal region $B\rightarrow X_{c}\ell\nu$ template normalizations are corrected before the fit and they are allowed to float separately within the correction factor uncertainties given in Table 3.
@@ V.5.13 $X_{u}$ hadronization modeling
An uncertainty on the fragmentation modeling of the $X_{u}$ system in non-resonant signal events is derived using the signal region $\pi^{\pm}$ multiplicity. It was checked beforehand in control regions that this variable is reasonably well modelled in $B\rightarrow X_{c}\ell\nu$ decays. In each of the three signal regions, the fit results are projected onto the $\pi^{\pm}$ multiplicity distribution and it is assumed that the remaining discrepancy between data and simulation is entirely due to the imperfect $X_{u}$ hadronization modeling. For each kinematic region, three scale factors are extracted in bins of the generator-level $\pi^{\pm}$ multiplicity of non-resonant $B\rightarrow X_{u}\ell\nu$ events (0, 1, 2 and more $\pi^{\pm}$) such that the simulated postfit $\pi^{\pm}$ multiplicity distribution matches perfectly the data distribution. The scale factors are applied to the pre-fit signal region distributions and the three fits are repeated. The difference between the measured $B\rightarrow X_{u}\ell\nu$ branching fraction and the branching fraction extracted from the rescaled distribution is taken as an uncertainty. Part of the uncertainty could originate from mismodeling of backgrounds in the $\pi^{\pm}$ multiplicity distribution. However, because of the relatively large signal purity in the three kinematic regions, the full difference is conservatively attributed to the $B\rightarrow X_{u}\ell\nu$ modelling.
@@ V.7 Prefit distributions
The three fitted variables in CR0,low and the signal region are shown in Figure 7. While the agreement between data and simulation appears to be relatively good in signal enriched regions, this is not the case in regions of the phase space where $B\rightarrow X_{c}\ell\nu$ events dominate. The shape mismodeling of this component is expected to be corrected by simultaneously fitting each distribution in the signal region and CR0,low. However, the normalization of the charm background in the signal regions appears to be overestimated after applying the normalization correction factors ($cf.$ Table 3). To evaluate the impact of the overestimated $B\rightarrow X_{c}\ell\nu$ normalizations, the injected signal setup described in Section V.6 is used. The injected signal normalization is fixed to unity and the $B\rightarrow X_{c}\ell\nu$ template in CRK,high is scaled up by an overestimation factor in the range [1.05, 1.10, 1.15, 1.20]. After correcting for the bias described in Section V.6, a slight bias increasing linearly with the $B\rightarrow X_{c}\ell\nu$ overestimation factor is observed. In the three signal regions considered, it is not possible to extract precisely the true $B\rightarrow X_{c}\ell\nu$ normalization without making assumptions about the $B\rightarrow X_{u}\ell\nu$ normalization, which is extracted from the fit. By checking the ratio of data to simulated data yields, it is estimated that the $B\rightarrow X_{c}\ell\nu$ normalization is overestimated by a factor roughly in the range 1.101.151.10-1.15 in all three regions. A flat correction factor is therefore extracted assuming that the overestimation factor is equal to 1.15. The entire difference between the measured branching fraction and the corrected one is added as an uncertainty. The $B\rightarrow X_{c}\ell\nu$ normalization bias correction is tested with different correction factors assuming various overestimation factors in the range 1.001.301.00-1.30. In all cases the assigned uncertainty is large enough to cover the true injected signal normalization after applying the bias correction.
@@ VI.1 Partial branching fractions
The postfit distributions in each of the three signal regions are shown in Figure 8. The excellent post-fit agreement is confirmed by the $p$-values obtained for each fit: 0.79, 0.20 and 0.84 for fits 1, 2 and 3, respectively. The event yields as extracted from fits 1, 2 and 3 are given in Table 5. In the kinematic regions considered for fits 1 and 2, the signal yields are similar to those reported in the latest Belle measurement [31] while the background yields are approximately 30–45% lower, despite using a data set that is about half the size of the 711 fb-1 data set collected by the Belle detector. The higher signal reconstruction efficiency can be explained by the improvements in the FEI performance compared to the Full Reconstruction algorithm used in Belle [53], most notably the inclusion of more $B$ decay chains. Furthermore, in this measurement, the continuum and $B\rightarrow X_{c}\ell\nu$ backgrounds are suppressed by relying on two neural networks which are optimized separately whereas a single BDT was used for that purpose in the Belle measurement. Finally, the increased reconstruction efficiency for $\pi_{s}$ helps to reject a larger fraction of $B\rightarrow D^{*}\ell\nu$ events.
@@
The total uncertainty breakdown in terms of individual sources of uncertainty is shown in Table 7 and the numerical impact of uncertainties related to the bias correction and the sample composition on the measured partial branching fractions is summarized in Table 8. These individual uncertainties are intended to provide insight into the sensitivity of the measurement to inputs and assumptions. However, it is not possible to account for correlations between fit parameters when calculating the individual uncertainties and the numbers quoted are therefore approximate. Hence, the sum in quadrature of all individual sources of uncertainty does not necessarily match the total uncertainty obtained from the fit. For all three measured branching fractions, the DFN model parameter uncertainty represents one of the largest sources of uncertainty. It is larger in fit 3 than in the other two fits because the $B\rightarrow X_{u}\ell\nu$ acceptance is the smallest among the three kinematic regions making the dependence on the model stronger. One of the leading uncertainties of the branching fractions measured with fits 2 and 3 is the hadronic system fragmentation modeling uncertainty. The result obtained from fit 2 is significantly impacted by the $B\rightarrow X_{c}\ell\nu$ composition uncertainty described in Section V.6 because the relative fractions of $B\rightarrow X_{c}\ell\nu$ sub-components are considerably different between the signal region and CRK,high when applying the $E_{\ell}^{B}$ and $M_{X}$ selections (see Table 10).
@@
Table 8: Breakdown of uncertainties added following the bias corrections described in Sections V.6 and V.7. In addition, we show the systematic, statistical and total uncertainties of the corrected branching fractions.
@@
The value of $|V_{ub}|$ can be determined from the measured partial branching fractions using Equation 2. We use for the value of the $B$ meson lifetime, the average of the $B^{+}$ and $B^{0}$ lifetimes, $\tau_{B}=1.578\pm 0.003$ ps [5]. Following Ref. [5], three different theoretical frameworks are used to extract $|V_{ub}|$:
@@
The values of $|V_{ub}|$ obtained from each model and each signal extraction fit using Equation 2 are summarized in Table 9. Theoretical uncertainties, $e.g.$, those due to the assumed value of the $b$ quark mass, impact both the determination of the branching fraction and the calculation of $|V_{ub}|$. However, the correlations between these quantities are difficult to determine due to the different parameterizations of non-perturbative effects used in the different theoretical calculations of $|V_{ub}|$ and in the DFN model used in our simulation. Therefore, similarly to what is done in the HFLAV averaging calculation, correlations between the branching fraction and partial decay rate theoretical uncertainties are ignored when extracting $|V_{ub}|$.
Table 9: Theoretical $B\rightarrow X_{u}\ell\nu$ decay rates $\Delta\tilde{\Gamma}$ (excluding the $|V_{ub}|^{2}$ term) and $|V_{ub}|$ obtained from three different theoretical predictions in three separate phase-space regions. The values of $\Delta\tilde{\Gamma}$ are used to compute $|V_{ub}|$ following Equation 2. The uncertainties on $|V_{ub}|$ are statistical, systematic and theoretical, respectively.

### v2

@@ V.1 Statistical model
The statistical model relies on the simultaneous measurement of event counts $n$ in two independent regions [^4] Referred to as channels in the PYHF documentation. – a signal region and a control region – using binned distributions of a given variable. These distributions are modeled as the sum of three different components (i.e. samples or templates, specified in Sec. V.3). Within the statistical model, a set of free and constrained parameters, $\bm{\eta}$ and $\bm{\chi}$ respectively, parametrize the variations of each template yield. The impact of each nuisance parameter $\chi$ is limited by a Gaussian constraint term $c_{\chi}$. The constraints are defined using a set of auxiliary measurements $\bm{a}$ (one measurement $a_{\chi}$ for each parameter), which correspond to component yield variations caused by various systematic effects. Thus, the statistical model $f$, in its most general form, can be written as
@@ V.2 Fitted regions and $B\rightarrow X_{c}\ell\nu$ shape correction
As detailed in Sec. IV.2, $B\rightarrow X_{c}\ell\nu$ decays appear to be poorly modeled. Normalization correction factors are calculated and applied before the fit whereas the correction of the $B\rightarrow X_{c}\ell\nu$ component shape is implemented directly into the fit. In order to constrain the shape of the $B\rightarrow X_{c}\ell\nu$ background in the signal region, distributions are fitted simultaneously in the signal region and CR0,low, the latter being orthogonal to the signal region and enriched in $B\rightarrow X_{c}\ell\nu$ events. This method not only corrects the $B\rightarrow X_{c}\ell\nu$ shape in the signal region, it also properly accounts for all correlations between correction factors and other fit parameters to determine a $B\rightarrow X_{u}\ell\nu$ branching fraction value that incorporates uncertainties related to the $B\rightarrow X_{c}\ell\nu$ background corrections. Using any control region to correct the shape of the $B\rightarrow X_{c}\ell\nu$ component relies on the assumption that the disagreement between data and simulation is similar in the signal region and the chosen control region. This assumption is supported by the fact that the ratio of $B\rightarrow X_{c}\ell\nu$ data over simulated data yields follows a similar trend in relevant kinematic variables in all three control regions as illustrated in Fig. 6. To choose the control region used for correcting the signal region mismodeling, two considerations must be taken into account: the similarity in composition and shape of the $B\rightarrow X_{c}\ell\nu$ component in the control and signal region and the sample size relative to the signal region. Based on these two considerations, CR0,low is chosen for the fit.
@@ V.3 Fit templates
  • Signal ($B\rightarrow X_{u}\ell\nu$): this template includes all signal events whose generated kinematics fall in the chosen phase-space region (signal-in). Signal decays are modeled via the DFN framework [30]. The template is assigned an unconstrained normalization factor in the statistical model whose postfit value can be directly translated into a measurement of the branching fraction. This factor is fixed to its value from simulation in CR0,low where its contribution is negligible. An additional template (signal-out) is defined to cover the small fraction of signal events which pass the reconstruction-level selections given in Table 2 but whose generated kinematic parameters fall outside the target signal region (based on the generator-level selections). The resolution of the $E_{\ell}^{B}$ distribution is extremely good, allowing the $B\rightarrow X_{u}\ell\nu$-out yield to be neglected in fit 1. In fits 2 and 3, this component is kept as it represents 2% and 5% of all signal events in their respective signal regions. The template is assigned the same unconstrained normalization as the signal-in template but nuisance parameters can vary its normalization and shape independently.
  • $B\rightarrow X_{c}\ell\nu$ background: this template represents the largest source of background. It is composed of $B\rightarrow D\ell\nu$, $B\rightarrow D^{*}\ell\nu$, $B\rightarrow D^{**}\ell\nu$ and $B\rightarrow D^{**}_{\textrm{Gap}}\ell\nu$ decays. To capture the shape of this template in CR0,low, one free-floating normalization factor per bin, common to the signal and control-region distributions, is used. The normalization of this template is corrected prior to fitting, following the procedure described in Sec. IV.3. A constrained normalization parameter is assigned to the $B\rightarrow X_{c}\ell\nu$ template in each region to float the overall normalization within the uncertainties given in Table 3.
@@ V.5.1 $B\rightarrow X_{u}\ell\nu$ modeling
The resonant and nonresonant branching fractions (given in Table 1) are varied within their respective uncertainties therefore yielding one nuisance parameter per decay type. The modeling of $B\rightarrow X_{u}\ell\nu$ resonant modes relies on assumed form factor models which depend on a set of parameters extracted from lattice QCD and dedicated measurements. The form factor parameter variations are calculated using the eFFORT [32] package. The parameter values, uncertainties and correlation matrices are taken from the references given for each model in Sec. II. Two nuisance parameters are added to represent the uncertainties in $m_{b}^{\mathrm{KN}}=4.66\pm 0.04$ GeV and $a^{\mathrm{KN}}=1.3\pm 0.5$, including their covariance. In addition, to factor in differences between inclusive $B\rightarrow X_{u}\ell\nu$ models, a set of events is simulated via the Bosch, Lange, Neubert, Paz (BLNP) framework [42, 23]. The difference between the nominal fitted distribution and the BLNP distribution is added as a symmetric uncertainty. Moreover, applying a kaon veto to suppress charm background in the signal region also rejects $B\rightarrow X_{u}\ell\nu$ events where the $X_{u}$ system hadronizes as a pair of kaons. The branching fraction of $B\to K\bar{K}(X)\ell\nu$ decays (where $X$ is a charmless hadronic system) is not known and an uncertainty related to this contribution is therefore extracted by varying the production of $s$ quarks relative to $u$ or $d$ production, $\gamma_{s}$. The two values measured by the TASSO [8] ($\gamma_{s}=0.35\pm 0.05$) and JADE [14] ($\gamma_{s}=0.27\pm 0.06$) collaborations are averaged and the uncertainty is chosen to cover the full spread of $1\sigma$ variations from the two values, giving $\gamma_{s}=0.30\pm 0.09$ (which corresponds to a branching fraction of roughly 0.02%). Three sets of events with $\gamma_{s}$ equal to 0.21, 0.30 and 0.39 are produced. The sample simulated with $\gamma_{s}$ equal to its central value is used as nominal and the variations with respect to the other two distributions are added as systematic uncertainties. For each variation related to the $B\rightarrow X_{u}\ell\nu$ modeling, the hybrid weights are recalculated (see Eq. 2).
@@ V.5.2 $B\rightarrow X_{c}\ell\nu$ modeling
All $B\rightarrow X_{c}\ell\nu$ and all $D$ meson decay branching fractions are varied within their respective uncertainties (see Table 1). The uncertainties of the unmeasured $B\to D^{(*)}\eta\ell\nu$ branching fractions are chosen to be equal to 100%. The branching fraction uncertainty of $B\to D_{0}^{*}\ell\nu$ decays is increased by about a factor of six to partially cover the large difference between the two measurements entering the HFLAV average calculation [12, 50]. The number of parameters related to the branching fractions of $D$ meson decays is truncated from 227 to one. For $B\rightarrow X_{c}\ell\nu$ decays, the form factor parameter variations are computed using the Hammer [15] package. The parameter values, uncertainties and correlation matrices are taken from the references given for each model in Sec. II.
@@ V.5.9 Continuum calibration
Simulated continuum events are corrected using the data-driven approach described in Sec. III.3. The full difference between the nominal distribution and the distribution corrected by the normalization factors is added as an uncertainty. Uncertainties on the continuum component shape also arise because of the limited size of the samples used to extract the correction factors. A total of 200 bootstrapped variations of the off-resonance experimental and simulated data samples are trained separately yielding as many sets of correction factors. The number of parameters implemented in the fit is truncated from 200 to five.
@@ V.5.11 Fit template normalizations
In fit 1, a normalization parameter is assigned to the other backgrounds template, which contributes to the total signal branching fraction uncertainty. The nuisance parameter is constrained by a Gaussian function with a width of 50% of the total component yields to vary the component normalization within a conservative range. In fits 2 and 3, the yields of nonsemileptonic events being negligible, this parameter is not added but other nuisance parameters related to this template can vary both its normalization and shape. Furthermore, the control region and signal region $B\rightarrow X_{c}\ell\nu$ template normalizations are corrected before the fit and they are allowed to float separately within the correction factor uncertainties given in Table 3.
@@ V.5.13 $X_{u}$ hadronization modeling
An uncertainty on the fragmentation modeling of the $X_{u}$ system in nonresonant signal events is derived using the signal region $\pi^{\pm}$ multiplicity. It was checked beforehand in control regions that this variable is reasonably well modeled in $B\rightarrow X_{c}\ell\nu$ decays. In each of the three signal regions, the fit results are projected onto the $\pi^{\pm}$ multiplicity distribution and it is assumed that the remaining discrepancy between data and simulation is entirely due to the imperfect $X_{u}$ hadronization modeling. For each kinematic region, three scale factors are extracted in bins of the generator-level $\pi^{\pm}$ multiplicity of nonresonant $B\rightarrow X_{u}\ell\nu$ events (0, 1, 2 and more $\pi^{\pm}$) such that the simulated postfit $\pi^{\pm}$ multiplicity distribution matches perfectly the data distribution. The scale factors are applied to the pre-fit signal region distributions and the three fits are repeated. The difference between the measured $B\rightarrow X_{u}\ell\nu$ branching fraction and the branching fraction extracted from the rescaled distribution is taken as an uncertainty. Part of the uncertainty could originate from mismodeling of backgrounds in the $\pi^{\pm}$ multiplicity distribution. However, because of the relatively large signal purity in the three kinematic regions, the full difference is conservatively attributed to the $B\rightarrow X_{u}\ell\nu$ modeling.
@@ V.7 Prefit distributions
The three fitted variables in CR0,low and the signal region are shown in Fig. 7. While the agreement between data and simulation appears to be relatively good in signal enriched regions, this is not the case in regions of the phase space where $B\rightarrow X_{c}\ell\nu$ events dominate. The shape mismodeling of this component is expected to be corrected by simultaneously fitting each distribution in the signal region and CR0,low. However, the normalization of the charm background in the signal regions appears to be overestimated after applying the normalization correction factors ($cf.$ Table 3). To evaluate the impact of the overestimated $B\rightarrow X_{c}\ell\nu$ normalizations, the injected signal setup described in Sec. V.6 is used. The injected signal normalization is fixed to unity and the $B\rightarrow X_{c}\ell\nu$ template in CRK,high is scaled up by an overestimation factor in the range [1.05, 1.10, 1.15, 1.20]. After correcting for the bias described in Sec. V.6, a slight bias increasing linearly with the $B\rightarrow X_{c}\ell\nu$ overestimation factor is observed. In the three signal regions considered, it is not possible to extract precisely the true $B\rightarrow X_{c}\ell\nu$ normalization without making assumptions about the $B\rightarrow X_{u}\ell\nu$ normalization, which is extracted from the fit. By checking the ratio of data to simulated data yields, it is estimated that the $B\rightarrow X_{c}\ell\nu$ normalization is overestimated by a factor roughly in the range 1.101.151.10-1.15 in all three regions. A flat correction factor is therefore extracted assuming that the overestimation factor is equal to 1.15. The entire difference between the measured branching fraction and the corrected one is added as an uncertainty. The $B\rightarrow X_{c}\ell\nu$ normalization bias correction is tested with different correction factors assuming various overestimation factors in the range 1.001.301.00-1.30. In all cases the assigned uncertainty is large enough to cover the true injected signal normalization after applying the bias correction.
@@ VI.1 Partial branching fractions
The postfit distributions in each of the three signal regions are shown in Fig. 8. The excellent post-fit agreement is confirmed by the $p$-values obtained for each fit: 0.79, 0.20 and 0.84 for fits 1, 2 and 3, respectively. The event yields as extracted from fits 1, 2 and 3 are given in Table 5. In the kinematic regions considered for fits 1 and 2, the signal yields are similar to those reported in the latest Belle measurement [27] while the background yields are approximately 30–45% lower, despite using a dataset that is about half the size of the 711 fb-1 dataset collected by the Belle detector. The higher signal reconstruction efficiency can be explained by the improvements in the FEI performance compared to the Full Reconstruction algorithm used in Belle [33], most notably the inclusion of more $B$ decay chains. Furthermore, in this measurement, the continuum and $B\rightarrow X_{c}\ell\nu$ backgrounds are suppressed by relying on two neural networks which are optimized separately whereas a single BDT was used for that purpose in the Belle measurement. Finally, the increased reconstruction efficiency for $\pi_{s}$ helps to reject a larger fraction of $B\rightarrow D^{*}\ell\nu$ events.
@@
The total uncertainty breakdown in terms of individual sources of uncertainty is shown in Table 7 and the numerical impact of uncertainties related to the bias correction and the sample composition on the measured partial branching fractions is summarized in Table 8. These individual uncertainties are intended to provide insight into the sensitivity of the measurement to inputs and assumptions. However, it is not possible to account for correlations between fit parameters when calculating the individual uncertainties and the numbers quoted are therefore approximate. Hence, the sum in quadrature of all individual sources of uncertainty does not necessarily match the total uncertainty obtained from the fit. For all three measured branching fractions, the DFN model parameter uncertainty represents one of the largest sources of uncertainty. It is larger in fit 3 than in the other two fits because the $B\rightarrow X_{u}\ell\nu$ acceptance is the smallest among the three kinematic regions making the dependence on the model stronger. One of the leading uncertainties of the branching fractions measured with fits 2 and 3 is the hadronic system fragmentation modeling uncertainty. The result obtained from fit 2 is significantly impacted by the $B\rightarrow X_{c}\ell\nu$ composition uncertainty described in Sec. V.6 because the relative fractions of $B\rightarrow X_{c}\ell\nu$ subcomponents are considerably different between the signal region and CRK,high when applying the $E_{\ell}^{B}$ and $M_{X}$ selections (see Table 10).
@@
Table 8: Breakdown of uncertainties added following the bias corrections described in Secs V.6 and V.7. In addition, we show the systematic, statistical and total uncertainties of the corrected branching fractions.
@@
The value of $|V_{ub}|$ can be determined from the measured partial branching fractions using Eq. 1. We use for the value of the $B$ meson lifetime, the average of the $B^{+}$ and $B^{0}$ lifetimes, $\tau_{B}=1.578\pm 0.003$ ps [13]. Following Ref. [13], three different theoretical frameworks are used to extract $|V_{ub}|$:
@@
The values of $|V_{ub}|$ obtained from each model and each signal extraction fit using Eq. 1 are summarized in Table 9. Theoretical uncertainties, $e.g.$, those due to the assumed value of the $b$ quark mass, impact both the determination of the branching fraction and the calculation of $|V_{ub}|$. However, the correlations between these quantities are difficult to determine due to the different parameterizations of nonperturbative effects used in the different theoretical calculations of $|V_{ub}|$ and in the DFN model used in our simulation. Therefore, similarly to what is done in the HFLAV averaging calculation, correlations between the branching fraction and partial decay rate theoretical uncertainties are ignored when extracting $|V_{ub}|$.
Table 9: Theoretical $B\rightarrow X_{u}\ell\nu$ decay rates $\Delta\tilde{\Gamma}$ (excluding the $|V_{ub}|^{2}$ term) and $|V_{ub}|$ obtained from three different theoretical predictions in three separate phase-space regions. The values of $\Delta\tilde{\Gamma}$ are used to compute $|V_{ub}|$ following Eq. 1. The uncertainties on $|V_{ub}|$ are statistical, systematic and theoretical, respectively.